A certain drug is made from only two ingredients: compound A and compound B. There are 5 milliliters of compound A used for every 8 milliliters of compound B. If a chemist wants to make 884 milliliters of the drug, how many milliliters of compound A are needed?

Answer by ankor@dixie-net.com(22740)

A certain drug is made from only two ingredients: compound A and compound B.
There are 5 milliliters of compound A used for every 8 milliliters of compound B.
If a chemist wants to make 884 milliliters of the drug, how many milliliters of compound A are needed?
:
We know the ratio of A to B is 5:8
let x = the multiplier
then
5x = amt of compound A
and
8x = amt of compound B
:
5x + 8x = 884
13x = 884
x = 884/13
x = 68 is the multiplier
therefore
5(68) = 340 millilters of compound A needed
:
:
Confirm our answer. Find the amt of B
8(68) = 544 ml of B
340 + 544 = 884